# A Survey on Causal Discovery: Theory and Practice

### v1

@@ 4.2 No Acyclicity Assumption
These additional patterns arise from a fundamental problem: the algorithms is not complete, and, therefore, there may be features common to all graphs in the same equivalence class that are not present in the output PAG (i.e. it is not the most informative PAG). While not being complete in the same sense of the previous algorithms, CDD is d-separation complete, meaning that the resulting PAG represents an equivalence class with a single graph, i.e. it encodes all the needed conditional independecies. Therefore, CCD is useful when one is interested in querying the resulting graph about dependencies, but lacks the capability to represent every causal edge by definition, in contrast to others algorithms. This limitation makes it less suitable for the definition of SCMs, especially when one is interested in the form of the functional set.
Linear Non-Gaussian (LiNG)
The LiNGAM algorithm can be adapted to the cyclic setting by weakening the acyclicity assumption. Specifically, instead of targeting a DAG, LiNG (or LiNG-D family) [48] try to recover a simple graph (i.e. without self-loops) by forcing all entries on the diagonal of the B matrix to be zero.
While LiNGAM output could be seen as a set of admissible models that contains a single model (i.e. the model is identifiable), the cyclic variant usually admits more than one causal graph at the time. In fact, the acyclicity assumption that allowed to find the row-permutation of B that best fits the given data set is missing. The authors then suggest to limit the discovery procedure to the $k$-th best assignment, following the intuition that permutations associated to inadmissible models would score poorly asymptotically. This approach selects one single model from the equivalent class (i.e. returning set).
@@
$\sigma$-Connection Graphs
From the concept of $\sigma$-separation, one can derive a MG where conditional independencies are expressed in the presence of cycles and latent variables, namely a $\sigma$-Connection Graph ($\sigma$-CG). An algorithm to learn this structures from data has been developed [24] as a natural extension of the work presented in [36]. The causal discovery problem is re-casted as a continuous optimization problem based on the following loss function:
$\mathcal{L}(G,\textbf{S})=\sum\lambda_{i}(\mathbbm{1}_{\lambda_{i}>0}-\mathbbm{1}_{X_{i}\perp\!\!\!\perp_{G}Y_{i}|\textbf{Z}_{i}})$ (4.2)
where S is a set of conditional independence statements expressed as $\textbf{S}=\big{(}(X_{i},Y_{i},\textbf{Z}_{i},\lambda_{i})\big{)}^{n}_{i=1}$, where $X_{i},Y_{i}$ and $\textbf{Z}_{i}$ are variables in V and $\lambda_{i}\in\mathbb{R}\cup\{-\infty,+\infty\}$ encodes the confidence of probabilistic conditional independence $X_{i}\perp\!\!\!\perp_{P}Y_{i}|\textbf{Z}_{i}$ as a constraint.
The $\lambda_{i}$ weights are evaluated using the indicator function $\mathbbm{1}$ to constrain the conditional dependence between variables. Therefore, Equation 4.2 quantify the amount of observations against the proposed causal graph based on the observed data. During the experimental evaluation, authors relied on the weights proposed in [53]:
@@
Minimizing the loss function may lead to multiple optimal solutions, where each solution $G$ is an instance of the actual equivalence class $[G]$. Indeed, as for d-separation and CPDAGs, the $\sigma$-separation criterion and the associated $\sigma$-CGs take into account possible undirected edges that are invariant for any causal graph belonging to the same equivalence class.
This algorithm has been benchmarked against synthetic data in low-dimensional setting. While the recovery metrics show consistent performances across the experiments, especially when increasing the number of interventions, it is clear that the main limitation of this approach is linked with the $\sigma$-separation encoding, as noted by [78]. Indeed, the separation checks are preformed using Answer Set Programming (ASP), a declarative logic programming language, which slows down the learning procedure.
bcause
@@
The bcause algorithm [77] explores the search space in a tree-like visit guided by an objective function that determines the weight of a potential solution. During the discovery phase, any edge of an intermediate result $G$ is either absent, present or undecided. Before the actual branching step, the lower bound of the given objective function for the current partial solution $G^{\prime}$ is computed. If such bound is higher than the weight obtained by the previous solution $G$, the branch can be closed and the algorithm backtracks. Otherwise, if $G^{\prime}$ contains at least one undecided edge, the procedure branches recursively in two directions: one in which said edge is set as present and the other marked as absent. Finally, if the branch cannot be closed and $G$ has no undecided edge, then the current solution $G^{\prime}$ is updated if and only if the evaluation of the objective function results in a lower weight. The search procedure will return $G$ as a globally optimal solution.
Since the causal discovery problem is inherently exponential, an exact search algorithm is unfeasible in the general setting. However, if both the objective function and its lower bound can be efficiently evaluated, a constrained space for a low dimensional problem can be effectively explored. For example, the authors benchmark their method under different conditions, showing that assuming acyclicity result in a lower execution time. Moreover, the algorithm maintains a set of constraints satisfied by the local solution and updates them incrementally. Therefore, any incompatible extension of the current solution is ruled out by leveraging a linear programming solver, reducing the total number of evaluation needed.
@@ 5 Causal Discovery with Interventions
The ladder of causation, also called the causal hierarchy, is an ordered structure of composed by three layers, where each layer is mapped to a cognition level: observational, interventional and counterfactual. A level inherently defines the set of potential queries that can be answered with the given information associated to it.
@@
Layer
Question
Method
Observational
How would seeing $X$ change my belief in $Y$?
Un/Supervised Learning
Interventional
What happens to $Y$ if I do $X$?
Reinforcement Learning
Counterfactual
What would have happened to $Y$ if I had done $X^{\prime}$ instead of $X$?
Structural Causal Model
@@
$do$-calculus
Queries that are usually expressed in natural language can be rephrased in terms of probability distribution by introducing the $do$ operator, whenever possible [^3] We restrict ourselves to a minimal introduction of the $do$-calculus, aiming to formally represent the set of concepts that are essential for the causal discovery scenario. For a broader discussion on identification and estimation of the causal effect, refer to [96].
@@
[…]
@@
[…]
@@
[…]
@@ 5.2 Types of Interventions
Indeed, $do$-calculus enables us to express perfect interventions in a operative framework, but there are other types of interventions that cannot be expressed using this notation.
@@
An intervention is said to be imperfect (or parametric, soft) [55] if it does not remove the causal dependence that affects the intervention target, but alters the functions that represents such dependence.
@@ 5.3 Defining the Intervention Target
[…]
@@
$\exists\textbf{I}:X\not\in\textbf{I}\in\mathcal{I},\quad\forall X\in\textbf{V}$ (5.4)
Essentially, a conservative family is a family that allows the existence of at least one intervention target which does not intervene on an specific variable. This property guarantees that there is at least one experiment in the family which does not alter the behaviour of such variable if performed.
In this settings, a conservative family allows to observe the influence of a (known) set of targets on at least one unaffected variable, enabling the possibility of disentangling such effect, especially when compared to the other experiments in the whole family.
@@
where $P^{(\varnothing)}$ is the observational distribution of the variables which were not included in the intervention target, if any.
@@
Two causal graphs $G$ and $H$ belong to the same interventional Markov equivalence class w.r.t. the intervention family $\mathcal{I}$ ($\mathcal{I}$-MEC) [40, 45] if they are $\mathcal{I}$-equivalent. As for the observational setting, the $\mathcal{I}$-MEC of a graph $G$, denoted by $[G]_{\mathcal{I}}$, represents the set of possible causal graphs that are interventionally equivalent.
@@
An edge $(X\rightarrow Y)$ in $G$ is $\mathcal{I}$-contradictory if as least one of the following conditions holds:
  • $\exists\textbf{S}\subset Ne(Y)\,\backslash\,\{X\}$ such that $\forall\textbf{I}\in\mathcal{I}_{X\backslash Y}$ we observe $P^{(\textbf{I})}(Y\,|\,\textbf{S})=P^{(\varnothing)}(Y\,|\,\textbf{S})$, or
  • $\forall\textbf{S}\subset Ne(X)\,\backslash\,\{Y\}$ such that $\exists\textbf{I}\in\mathcal{I}_{Y\backslash X}$ we observe $P^{(\textbf{I})}(X\,|\,\textbf{S})\neq P^{(\varnothing)}(X\,|\,\textbf{S})$.
@@ 5.4 Learning with Interventions
Sometimes researchers want to observe the effect of an intervention on one single variable at the time, but there are settings in which this is not possible or it is inconvenient. Therefore, multi-variable interventions must be addressed as a special case of a generic intervention target.
@@
When each intervention target contains a single variable at the time, the number of experiments needed to collect enough evidence to identify the causal graph is $n-1$, with $n$ the number of variables [23]. Indeed, if one intervention would enable the identification of the causal edges incoming into the only variable contained in the intervention target, then the $n$-th intervention would be redundant.
In the case of intervention targets with more then one variable, only $\lfloor\log(n)\rfloor+1$ interventions are necessary and sufficient in the worst case scenario [23], where the causal graph is the complete graph. Since this worst case is improbable, $O(\log\log(n))$ can be achieved as lower bound with high probability in the multi-variable setting with a randomized intervention scheme [35], that is, it is possible to plan the experimental design in advance to minimize the number of interventions.
@@
An other problem that one may face during structural learning with interventional data is the uncertainty related to the interventional targets. There are scenarios in which it is known that an intervention has been performed, but it is unclear which is the exact set of variables that has been affected by such intervention. In this case, an additional layer of complexity is added in order to properly handle the less informative setting of unknown intervention targets.
@@ 5.5 Interventional Algorithms
A major contribution of this work is related to formalization of the interventional setting. Indeed, while the algorithm itself does not differ significantly from the observational one in terms of overall design, the performance improvements are relevant, as expected by transitioning from the first to the second layer of the causal hierarchy. This is an interesting example of how observational techniques can be adapted to the interventional setting with ease, once the theoretical aspects of both the intervention distribution and the intervention targets are addressed properly.
Interventional Greedy Permutation (IGSP)
@@
Formally, let $\rho$ be a permutation of vertices of a causal graph $G$. The space on which such permutation lays is a polytope called permutahedron. A possible representation of this mathematical object is indeed another graph, where each vertex corresponds to a permutation $\rho$ and each edge between two permutations encodes a transposition of the vertices. The goal of a permutation-base causal discovery algorithm is to find a permutation $\rho^{*}$, consistent with the topological order of the true causal graph $G^{*}$, that optimizes a given score function. The search procedure traverses the permutahedron using a depth-first approach starting from an initial permutation $\rho$. For each permutation $\tau$ visited, if $G_{\tau}$ yields a better score than $G_{\rho}$ then $\rho$ is set to $\tau$. The traversal is restarted from the updated $\rho$, until no such $\tau$ is found.
@@
Joint Causal Inference with FCI (FCI-JCI)
@@
In this framework, the set of observed variables is split into two disjoint sets: system variables X and context variables C. While the former set contains the variables that have been observed during an experiment, the latter set describes under which conditions such system has been observed, following the classical distinction between endogenous and exogenous variables respectively.
@@
The generic JCI procedure can adapted to any observational causal discovery algorithm by following four steps: (i) add the context variables, (ii) pool data together by setting the values of the context variables, (iii) address faithfulness violations between contexts, if any, (iv) execute the selected observational learning algorithm. Authors provide reference adaptations for multiple algorithms, such as FCI.
The FCI-JCI variant is particularly of interest, provided that it inherits the strength points of FCI in the causally insufficient setting. Various combinations of the three assumptions were tested, showing that FCI123 (i.e. all three assumptions made) is less accurate in general, but significantly faster than others solutions, allowing its application in more complex scenarios with a sensible number of variables.
Unknown Intervention Targets using $\Psi$-FCI
@@
The key idea is that a pair of intervention targets $\textbf{I},\textbf{J}\in\mathcal{I}$ can be used to identify a unique interventional mechanism that encompasses both targets. Let $G$ be a causal graph and $\mathcal{I}$ an intervention family. The induced set of interventional probability distributions $P^{(\mathcal{I})}=\{P^{(\textbf{I}_{0})},P^{(\textbf{I}_{1})},\dots,P^{(\textbf{I}_{n})}\}$ satisfies the $\Psi$-Markov property if the following holds true for any disjoint subsets of variables Y, Z and W:
@@
[…]
@@
    if $\big{(}\textbf{Y}\perp\!\!\!\perp\textbf{K}\,|\,(\textbf{W}\,\backslash\,\textbf{W}_{\textbf{K}})\big{)}$ holds true in $G_{\underline{\textbf{W}_{\textbf{K}}}\overline{\textbf{R}(\textbf{W})}}$ for all $\textbf{I},\textbf{J}\in\mathcal{I}$, where K is the symmetric difference of I and J, $=\textbf{W}\cap\textbf{K}$ W ∩ K, $\textbf{R}=\textbf{K}\,\backslash\,\textbf{W}_{\textbf{K}}$ and finally $\textbf{R}(\textbf{W})=\textbf{R}\,\backslash\,An(\textbf{W})$ w.r.t. $G$.
While Equation 5.10 is essentially derived from observational Markov equivalence, Equation 5.11 is related to the distributional invariances across pairs if intervention targets w.r.t. the associated intervention graph. Indeed, if I and J are the true intervention targets for $P^{(\textbf{I})}$ and $P^{(\textbf{J})}$, they must satisfy the invariance for the interventional distributions when separation holds in a given intervention graph.
@@
The authors then recast the augmented graph proposed by [11, 25], adding a set of utility vertices that are analogous to the context vertices proposed by [61]. Therefore, the output of the former can be compared to the latter using the related augmented graph, showing that the accuracy of the edge orientations recovered by their $\Psi$-FCI variant is superior then the one proposed by FCI-JCI.
backShift
@@
The key idea of this technique is to represent the the target SCM $M$ as:
[…]
@@
$\widetilde{\textbf{D}}=\operatorname*{\arg\!\min}_{\textbf{D}^{\prime}\in\mathcal{D}}\sum_{\textbf{I}\in\mathcal{I}}\mathcal{L}(\textbf{D}^{\prime}\Delta\boldsymbol{\Sigma}^{(\textbf{I})}\textbf{D}^{\prime T})$ (5.13)
[…]
@@
Authors compare their solution to the LiNG observational alternative, taking advantage of the interventional asymmetries arising from the additional information contained in the data. The results show that backShift is capable of dealing with both interventions and latent variables under mild assumptions, outperforming LiNG in both observational and interventional setting. Moreover, the computational complexity is $O(|\mathcal{I}|\cdot n^{2}\cdot m)$, with $n$ representing the number of the variables and $m$ representing the sample size, which allows its application in high-dimensional settings.
bcause+
@@
Differentiable Causal Discovery with Interventions (DCDI)

### v2

@@ 4.2.1 Cyclic Causal Discovery (CCD)
These additional patterns arise from a fundamental problem: the algorithms is not complete (Definition 2.12), and, therefore, there may be features common to all graphs in the same equivalence class that are not present in the output PAG (i.e. it is not the most informative PAG). While not being complete in the same sense of the previous algorithms, CCD is d-separation complete, meaning that the resulting PAG represents an equivalence class with a single graph, i.e. it encodes all the needed conditional independencies. Therefore, CCD is useful when one is interested in querying the resulting graph about dependencies, but lacks the capability to represent every causal edge by definition, in contrast to others algorithms. This limitation makes CCD less suitable for the definition of SCMs, especially when one is interested in the form of the functional set.
4.2.2 Linear Non-Gaussian (LiNG)
The LiNGAM algorithm can be adapted to the cyclic setting by weakening the acyclicity assumption. Specifically, instead of targeting a DAG, LiNG (or LiNG-D family) [49] tries to recover a simple graph (i.e. without self-loops) by forcing all entries on the diagonal of the B matrix to be zero.
While LiNGAM output could be seen as a set of admissible models that contains a single model (i.e. the model is identifiable), the cyclic variant usually admits more than one causal graph at a time. In fact, the acyclicity assumption that allowed to find the row-permutation of B that best fits the given dataset is missing. The authors then suggest to limit the discovery procedure to the $k$-th best assignment, following the intuition that permutations associated to inadmissible models would score poorly asymptotically. This approach selects one single model from the equivalent class (i.e. returning set).
@@ 4.2.2 Linear Non-Gaussian (LiNG)
4.2.3 $\sigma$-Connection Graphs
From the concept of $\sigma$-separation, one can derive a MG where conditional independencies are expressed in the presence of cycles and latent variables, namely a $\sigma$-Connection Graph ($\sigma$-CG). An algorithm to learn this structure from data has been developed [52] as a natural extension of the work presented in [85]. The causal discovery problem is re-casted as a continuous optimization problem based on the following loss function:
$\mathcal{L}(G,\textbf{S})=\sum_{i=1}^{n}\lambda_{i}(\mathbbm{1}_{\lambda_{i}>0}-\mathbbm{1}_{X_{i}\perp\!\!\!\perp_{G}Y_{i}|\textbf{Z}_{i}})$ (4.2)
where S is a set of conditional independence statements expressed as $\textbf{S}=\big{(}(X_{i},Y_{i},\textbf{Z}_{i},\lambda_{i})\big{)}^{n}_{i=1}$, $X_{i},Y_{i}$ and $\textbf{Z}_{i}$ are variables in V, $\lambda_{i}\in\mathbb{R}\cup\{-\infty,+\infty\}$ encodes the confidence of probabilistic conditional independence $X_{i}\perp\!\!\!\perp_{P}Y_{i}|\textbf{Z}_{i}$ as a constraint and $\mathbbm{1}$ is the indicator function which assumes the value one when the constraint is satisfied.
The $\lambda_{i}$ weights are evaluated using the indicator function $\mathbbm{1}$ to constrain the conditional dependence between variables. Therefore, Equation 4.2 quantifies the observations against the proposed causal graph based on the observed data. During the experimental evaluation, authors relied on weights proposed in [86]:
@@ 4.2.3 $\sigma$-Connection Graphs
Minimizing the loss function may lead to multiple optimal solutions, where each solution $G$ is an instance of the actual equivalence class $[G]$. Indeed, as for d-separation and CPDAGs, the $\sigma$-separation criterion and the associated $\sigma$-CGs take into account possible undirected edges that are invariant for any causal graph belonging to the same equivalence class $[G]$.
This algorithm has been benchmarked against synthetic data in a low-dimensional setting. While the recovery metrics show consistent performance across the experiments, especially when increasing the number of interventions, it is clear that the main limitation of this approach is linked with the $\sigma$-separation encoding, as noted in [57]. Indeed, the separation checks are performed using Answer Set Programming (ASP), a declarative logic programming language, which slows down the learning procedure.
4.2.4 bcause
@@
The bcause algorithm [51] explores the search space in a tree-like visit guided by an objective function that determines the weight of a potential solution. During the discovery phase, any edge of an intermediate result $G$ is either absent, present or undecided. Before the actual branching step, the lower bound of the given objective function for the current partial solution $G^{\prime}$ is computed. If such bound is higher than the weight obtained by the previous solution $G$, the branch can be closed and the algorithm backtracks. Otherwise, if $G^{\prime}$ contains at least one undecided edge, the procedure branches recursively in two directions: one in which said edge is set as present and the other marked as absent. Finally, if the branch cannot be closed and $G$ has no undecided edges, then the current solution $G^{\prime}$ is updated if and only if the evaluation of the objective function results in a lower weight. The search procedure will return $G$ as a globally optimal solution.
Since the causal discovery problem is inherently exponential, an exact search algorithm is unfeasible in the general setting. However, if both the objective function and its lower bound can be efficiently evaluated, a constrained space for a low dimensional problem can be effectively explored. For example, the authors benchmark their method under different conditions, showing that assuming acyclicity results in a lower execution time. Moreover, the algorithm maintains a set of constraints satisfied by the local solution and updates them incrementally. Therefore, any incompatible extension of the current solution is ruled out by leveraging a linear programming solver, reducing the total number of evaluations needed.
@@ 5 Causal Discovery with Interventions
The ladder of causation, also called the causal hierarchy, is an ordered structure composed by three layers, where each layer is mapped to a cognition level: observational, interventional and counterfactual. A level inherently defines the set of potential queries that can be answered with the given information associated to it.
@@
Layer Question Method
Observational How would seeing $X$ change my belief in $Y$? Un/Supervised Learning
Interventional What happens to $Y$ if I do $X$? Reinforcement Learning
Counterfactual What would have happened to $Y$ if I had done $X^{\prime}$ instead of $X$? Structural Causal Model
@@
dodo-calculus
Queries that are usually expressed in natural language can be rephrased in terms of probability distribution by introducing the $do$ operator, whenever possible [^4] We restrict ourselves to a minimal introduction of the $do$-calculus, aiming to formally represent the set of concepts that are essential for the causal discovery scenario. For a broader discussion on identification and estimation of the causal effect, refer to [88].. The $do$ operator represents an intervention on a given variable, e.g. $do(X=x)$ sets the value of $X$ to $x$. This notation is usually overloaded by extending the operator over sets of variables, e.g. $do(\textbf{X}=\textbf{x})$ with x a vector of values. Finally, the x vector can be omitted entirely for brevity.
@@
    if $(\textbf{Y}\perp\!\!\!\perp\textbf{Z}\,|\,\textbf{X},\textbf{W})$ holds true in $G_{\overline{\textbf{X}}}$,
@@
    if $(\textbf{Y}\perp\!\!\!\perp\textbf{Z}\,|\,\textbf{X},\textbf{W})$ holds true in $G_{\overline{\textbf{X}},\underline{\textbf{Z}}}$,
@@
    if $(\textbf{Y}\perp\!\!\!\perp\textbf{Z}\,|\,\textbf{X},\textbf{W})$ holds true in $G_{\overline{\textbf{X},\textbf{Z}(\textbf{W})}}$,
@@ 5.2 Types of Interventions
Indeed, $do$-calculus enables us to express perfect interventions in an operative framework, but there are other types of interventions that cannot be expressed using this notation.
@@
An intervention is said to be imperfect (or parametric, soft) [89] if it does not remove the causal dependence that affects the intervention target, but alters the functions that represent such dependence.
@@ 5.3 Defining the Intervention Target
Let $G$ be a causal graph. A subset $\textbf{I}\subset\textbf{V}$ is said to be an intervention target if it contains all and only the variables associated to an intervention over $G$.
@@
$\exists\textbf{I}\in\mathcal{I}:X\not\in\textbf{I},\quad\forall X\in\textbf{V}$ (5.4)
Essentially, a conservative family is a family that allows the existence of at least one intervention target that does not intervene on a specific variable. This property guarantees that there is at least one experiment in the family that does not alter the behaviour of such variable if performed.
In this setting, a conservative family allows to observe the influence of a (known) set of targets on at least one unaffected variable, enabling the possibility of disentangling such effect, especially when compared to the other experiments in the whole family.
@@
where $P^{(\varnothing)}$ is the observational distribution of the variables that were not included in the intervention target, if any.
@@
Two causal graphs $G$ and $H$ belong to the same interventional Markov equivalence class w.r.t. the intervention family $\mathcal{I}$ ($\mathcal{I}$-MEC) [55, 91] if they are $\mathcal{I}$-equivalent. As for the observational setting, the $\mathcal{I}$-MEC of a graph $G$, denoted by $G]_{\mathcal{I}}$, represents the set of possible causal graphs that are interventionally equivalent.
@@
An edge $(X\rightarrow Y)$ in $G$ is $\mathcal{I}$-contradictory if at least one of the following conditions holds:
[…]
@@ 5.4 Learning with Interventions
Sometimes researchers want to observe the effect of an intervention on one single variable at a time, but there are settings in which this is not possible or it is inconvenient. Therefore, multi-variable interventions must be addressed as a special case of a generic intervention target.
@@
When each intervention target contains a single variable at a time, the number of experiments needed to collect enough evidence to identify the causal graph is n1n-1, with $n$ the number of variables [92]. Indeed, if one intervention would enable the identification of the causal edges incoming into the only variable contained in the intervention target, then the $n$-th intervention would be redundant.
In the case of intervention targets with more than one variable, only $\lfloor\log(n)\rfloor+1$ interventions [^5] Where $\lfloor x\rfloor$ denotes the floor function that maps a real number $x$ to the greatest integer less than or equal to $x$. are necessary and sufficient in the worst case scenario [92], where the causal graph is the complete graph. Since this worst case is improbable, $O(\log\log(n))$ can be achieved as lower bound with high probability in the multi-variable setting with a randomized intervention scheme [93], that is, it is possible to plan the experimental design in advance to minimize the number of interventions.
@@
An other problem that one may face during structural learning with interventional data is the uncertainty related to the interventional targets [54, 55, 56]. There are scenarios in which it is known that an intervention has been performed, but it is unclear which is the exact set of variables that has been affected by such intervention. In this case, an additional layer of complexity is added in order to properly handle the less informative setting of unknown intervention targets.
@@ 5.5.1 Interventional GES (GIES)
A major contribution of this work is related to the formalization of the interventional setting. Indeed, while the algorithm itself does not differ significantly from the observational one in terms of overall design, the performance improvements are relevant, as expected by transitioning from the first to the second layer of the causal hierarchy. This is an interesting example of how observational techniques can be adapted to the interventional setting with ease, once the theoretical aspects of both the intervention distribution and the intervention targets are addressed properly.
5.5.2 Interventional Greedy Permutation (IGSP)
@@
Formally, let $\rho$ be a permutation of vertices of a causal graph $G$. The space on which such permutation lays is a polytope called permutahedron. A possible representation of this mathematical object is indeed another graph, where each vertex corresponds to a permutation $\rho$ and each edge between two permutations encodes a transposition of the vertices. The goal of a permutation-based causal discovery algorithm is to find a permutation $\rho^{*}$, consistent with the topological order of the true causal graph $G^{*}$, that optimizes a given score function. The search procedure traverses the permutahedron using a depth-first approach starting from an initial permutation $\rho$. For each permutation $\tau$ visited, if $G_{\tau}$ yields a better score than $G_{\rho}$ then $\rho$ is set to $\tau$. The traversal is restarted from the updated $\rho$, until no such $\tau$ is found.
@@
5.5.3 Joint Causal Inference with FCI (FCI-JCI)
@@
In this framework, the set of observed variables is split into two disjoint sets: system variables X and context variables C. While the former set contains the variables that have been observed during an experiment, the latter set describes under which conditions such system has been observed, following the classical distinction between endogenous and exogenous variables, respectively.
@@
The generic JCI procedure can be adapted to any observational causal discovery algorithm by following four steps: i) add the context variables, ii) pool data together by setting the values of the context variables, iii) address faithfulness violations between contexts, if any, iv) execute the selected observational learning algorithm. Authors provide reference adaptations for multiple algorithms, such as FCI.
The FCI-JCI variant is particularly of interest, provided that it inherits the strong points of FCI in the causally insufficient setting. Various combinations of the three assumptions were tested, showing that FCI123 (i.e. all three assumptions made) is less accurate in general, but significantly faster than others solutions, allowing its application in more complex scenarios with a sensible number of variables.
5.5.4 Unknown Intervention Targets using $\Psi$-FCI
@@
The key idea is that a pair of intervention targets $\textbf{I},\textbf{J}\in\mathcal{I}$ can be used to identify a unique interventional mechanism that encompasses both targets. Let $G$ be a causal graph and $\mathcal{I}$ an intervention family. The induced set of interventional probability distributions $P^{(\mathcal{I})}=\{P^{(\textbf{I}_{0})},P^{(\textbf{I}_{1})},\dots,P^{(\textbf{I}_{n})}\}$ satisfies the $\Psi$-Markov property if the following holds true for any Y, Z and W disjoint subsets of variables of V:
@@
    if $(\textbf{Y}\perp\!\!\!\perp\textbf{Z}\,|\,\textbf{W})$ holds true in $G$ for all $\textbf{I}\in\mathcal{I}$,
@@
[…]
While Equation 5.10 is essentially derived from observational Markov equivalence, Equation 5.11 is related to the distributional invariances across pairs of intervention targets w.r.t. the associated intervention graph. Indeed, if I and J are the true intervention targets for $P^{(\textbf{I})}$ and $P^{(\textbf{J})}$, they must satisfy the invariance for the interventional distributions when separation holds in a given intervention graph.
@@
The authors then recast the augmented graph proposed by [24, 33], adding a set of utility vertices that are analogous to the context vertices proposed by [28]. Therefore, the output of the former can be compared to the latter using the related augmented graph, showing that the accuracy of the edge orientations recovered by their $\Psi$-FCI variant is superior than the one proposed by FCI-JCI.
5.5.5 backShift
@@
The key idea of this technique is to represent the target SCM $M$ as:
$(\textbf{I}-\textbf{B})\textbf{x}=\textbf{c}+\textbf{e}$ (5.12)
@@
$\widetilde{\textbf{D}}=\operatorname*{\arg\!\min}_{\textbf{D}\in\mathcal{D}}\sum_{\textbf{I}\in\mathcal{I}}\mathcal{L}(\textbf{D}\Delta\boldsymbol{\Sigma}^{(\textbf{I})}\textbf{D}^{T})$ (5.13)
where $\textbf{D}=\textbf{I}-\textbf{B}$, $\mathcal{L}$ is the sum-of-squared loss function, $\mathcal{I}$ the family of targets.
@@
Authors compare their solution to the LiNG observational alternative, taking advantage of the interventional asymmetries arising from the additional information contained in the data. The results show that backShift is capable of dealing with both interventions and latent variables under mild assumptions, outperforming LiNG in both the observational and interventional settings. Moreover, the computational complexity is $O(|\mathcal{I}|\cdot n^{2}\cdot m)$, with $n$ representing the number of the variables and $m$ representing the sample size, which allows its application in high-dimensional settings.
5.5.6 bcause+
@@
5.5.7 Differentiable Causal Discovery with Interventions (DCDI)